OpenStudy (anonymous):

bryan wanted to increase the concentration of a liter of hydrochloric acid from 30% to 35%. How much pure acid should he add

OpenStudy (anonymous):

Is there a quantity it begins with?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it says a liter of hydrochloric acid

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

how much pure acid should he add

OpenStudy (anonymous):

idk the only way to get from 30 to 35 would be to add 5%

OpenStudy (anonymous):

Start with a table

OpenStudy (anonymous):

thats all i would think u would have to create a chart

OpenStudy (anonymous):

yea i did a table but its not makin since

OpenStudy (amistre64):

system of equations

OpenStudy (amistre64):

let h = 1 liter, and n be the amount to add in h+n = (1+n) .30h + 1n = .35(1+n)

OpenStudy (amistre64):

something similar to that

OpenStudy (anonymous):

Starting solution Total Amount: 1 liter %HCL acid: 30 Total Amount of HCL acid: .3 liter Solution Added: x liter %HCL acid : 100 Total amount of HCL acid: 1x Resulting Solution: 1+x %HCL acid: 35 Total amount of HCL: .35(1+x) .35(1+x)= x+.3

OpenStudy (amistre64):

seems we concur

OpenStudy (anonymous):

Yeah @amistre64 then just solve for x and plug it into the table

OpenStudy (amistre64):

x is the solution :)

OpenStudy (anonymous):

its 5%

OpenStudy (amistre64):

more like about .077 liters

OpenStudy (amistre64):

not sure how you would add 5% pure acid as a solution

OpenStudy (anonymous):

I think she's thinking a 30 % solution + 5% soultion would equal a 35% solution when Mixture problems work differently

OpenStudy (anonymous):

yea lol

OpenStudy (amistre64):

1liter at 30% + .077 liters at 100% = 1.077 liters at 35%

OpenStudy (anonymous):

The quantity of the starting solution has to be acknowledged which is in this case a liter

OpenStudy (anonymous):

The chemist is adding pure HCL which is .77 liters at 100% to make a 30 percent solution equal 35 percent

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=.30+%2B+n+%3D+.35%281%2Bn%29 .077 is what im getting, so im assuming a typo

OpenStudy (anonymous):

lol theirs no typo

OpenStudy (anonymous):

Yeah lol I didn't solve it I was just parroting your answer. @cali16 Lol he was talking to me.

OpenStudy (anonymous):

Yeah it's .077

OpenStudy (amistre64):

good luck yall ;)

OpenStudy (anonymous):

But do you get it @cali16

OpenStudy (anonymous):

yes now i do thx!!!

OpenStudy (anonymous):

No problem.